OpenStudy (anonymous):

A 6-foot spruce tree is planted 15 feet from a lighted streetlight whose lamp is 18 feet above the ground. How many feet long is the shadow of that tree?

jimthompson5910 (jim_thompson5910):

draw the picture to get 

jimthompson5910 (jim_thompson5910):

the two triangles are similar, so you can set up the proportion 18/6 = (15+x)/x and cross multiply to get 18x = 6(15+x) from here, you solve for x to get the length of the tree's shadow